In a certain plant, a dominant allele for one gene
Answer #1

The map distance is 22 cM can be calculated from the frequency of cross overs.

Here , Total gametes=381,

Yellow & Round=3, White & oblong=295, Yellow &oblong= 73, White & Round = 10.

We know, first two are normal gametes (as they are dominant and recessive) while the rest two are cross over gametes.

Therefore,

Frequency of the cross over gametes= number of crossover gametes divided by total number of gametes= 83/381 = 22 cM

Same way for semi-fertile,

Total gamete= 369

Frequency of crossover gamete= 82/369 = 22 cM